# PsychEval: A Multi-Session and Multi-Therapy Benchmark for High-Realism AI Psychological Counselor

### v2

@@ Abstract
To develop a reliable AI for psychological assessment, we introduce PsychEval, a multi-session, multi-therapy, and highly realistic benchmark designed to address three key challenges: 1) Can we train a highly realistic AI counselor? Realistic counseling is a longitudinal task requiring sustained memory and dynamic goal tracking. We propose a multi-session benchmark (spanning 6-10 sessions across three distinct stages) that demands critical capabilities such as memory continuity, adaptive reasoning, and longitudinal planning. The dataset is annotated with extensive professional skills, comprising over 677 meta-skills and 4577 atomic skills. 2) How to train a multi-therapy AI counselor? While existing models often focus on a single therapy, complex cases frequently require flexible strategies among various therapies. We construct a diverse dataset covering five therapeutic modalities alongside an integrative therapy with a unified three-stage clinical framework across six core psychological topics. 3) How to systematically evaluate an AI counselor? We establish a holistic evaluation framework with 18 therapy-specific and therapy-shared metrics across Client-Level and Counselor-Level dimensions. To We also construct over 2,000 diverse client profiles. Extensive experimental analysis fully validates the superior quality and clinical fidelity of our dataset. Crucially, PsychEval transcends static benchmarking to serve as a high-fidelity reinforcement learning environment that enables the self-evolutionary training of clinically responsible and adaptive AI counselors.
@@ 6.2 Comparison with Existing Benchmark
To contextualize PsychEval’s contribution, we compare it multidimensionally against representative datasets. As shown in Tables 4 and 2, PsychEval shifts focus from large-scale but shallow interactions to high-fidelity, longitudinal, and skill-aware assessment.
@@ Appendix D Prompts
D.1 Prompt for Client Profile
@@
D.2 Prompt for Therapeutic Plan
@@
D.3 Prompt for Dialogue Features
@@
D.4 Prompt for Session Objective
@@
D.5 Prompt for Skill Suggestion
@@
D.6 Prompt for Dialogue Generation
@@
D.7 Prompt for Client Memory Extraction
@@
D.8 Prompt for Client Merge
@@
D.9 Prompt for Dialogue Summary

### v3

@@ Abstract
To develop a reliable AI for psychological assessment, we introduce PsychEval, a multi-session, multi-therapy, and highly realistic benchmark designed to address three key challenges: 1) Can we train a highly realistic AI counselor? Realistic counseling is a longitudinal task requiring sustained memory and dynamic goal tracking. We propose a multi-session benchmark (spanning 6-10 sessions across three distinct stages) that demands critical capabilities such as memory continuity, adaptive reasoning, and longitudinal planning. The dataset is annotated with extensive professional skills, comprising over 677 meta-skills and 4577 atomic skills. 2) How to train a multi-therapy AI counselor? While existing models often focus on a single therapy, complex cases frequently require flexible strategies among various therapies. We construct a diverse dataset covering five therapeutic modalities alongside an integrative therapy with a unified three-stage clinical framework across six core psychological topics. 3) How to systematically evaluate an AI counselor? We establish a holistic evaluation framework with 18 therapy-specific and therapy-shared metrics across Client-Level and Counselor-Level dimensions. To We also construct over 2,000 diverse client profiles. Extensive experimental analysis fully validates the superior quality and clinical fidelity of our dataset. Crucially, PsychEval transcends static benchmarking to serve as a high-fidelity reinforcement learning environment that enables the self-evolutionary training of clinically responsible and adaptive AI counselors.Our datasets and evaluation framework are publicly available at https://github.com/ECNU-ICALK/PsychEval.
@@ 6.2 Comparison with Existing Benchmark
To contextualize PsychEval’s contribution, we compare it multidimensionally against representative datasets. As shown in Tables 2 and 4, PsychEval shifts focus from large-scale but shallow interactions to high-fidelity, longitudinal, and skill-aware assessment.
@@ 7 Conclusions and Future Work
8 Limitations
While PsychEval establishes a foundational benchmark for evaluating multi-session, multi-therapy AI psychological counseling, certain aspects reflect natural constraints of the current research landscape rather than shortcomings of the framework itself. First, the benchmark is built exclusively on textual interactions, consistent with the dominant paradigm in contemporary NLP. Although real-world counseling also relies on non-verbal cues such as vocal prosody, facial expressions, and body language, these modalities remain challenging to incorporate at scale and are beyond the scope of current text-based evaluation protocols. Second, the 369 case reports, curated from professional journals, emphasize clinically representative and ethically shareable trajectories, which necessarily underrepresent extreme scenarios such as acute suicidality or highly resistant clients. This reflects broader data availability challenges in mental health AI rather than a design flaw. Third, PsychEval is currently grounded in the Chinese linguistic and socio-cultural context, acknowledging that mental health expressions, help-seeking behaviors, and therapeutic norms vary across cultures, a dimension ripe for future expansion rather than a limitation of the current instantiation. Collectively, these points highlight promising avenues for future work, including multimodal integration, high-risk scenario modeling, and cross-cultural adaptation, while underscoring that PsychEval already provides a rigorous, clinically informed foundation for advancing responsible AI in mental health.
9 Ethical Statement
The PsychEval benchmark is constructed exclusively from 369 publicly available, peer-reviewed clinical case reports published in authoritative academic journals such as Theory and Practice of Psychological Counseling and Psychologist. These reports have already undergone formal ethical review and obtained informed consent in their original publication contexts, ensuring full compliance with clinical and research ethics standards. The dataset is derived solely from published scholarly sources and not from social media, private records, or unregulated platforms, and therefore entails no additional ethical risks related to data collection. Furthermore, PsychEval is released strictly for academic research under a responsible usage agreement, with clear disclaimers that AI systems trained on it must serve only as auxiliary tools or simulation environments and never as substitutes for human clinicians.
@@ Appendix C Statistic Information
Appendix D Experimental Settings and Dataset Construction Details
D.1 Model Configurations
For generating clinical dialogues, we employed GPT-5 with high reasoning effort to accurately follow complex clinical psychology instructions while preserving professional terminology and capturing the natural linguistic variability of clients. To balance evaluation effectiveness with computational cost, we selected DeepSeek-V3.1 as our primary evaluation model (LLM-as-a-Judge), owing to its strong performance in Chinese linguistic comprehension and logical reasoning. Due to budget constraints, we uniformly sampled 1,000 instances from each existing benchmark for comparative analysis.
D.2 Data Reconstruction and Quality Control
We implemented a rigorous, multi-stage quality control process to transform static clinical case reports into high-fidelity therapeutic dialogues. First, a timeline constraint was algorithmically enforced during generation to ensure the AI counselor only accessed information disclosed up to the current session, eliminating any “anticipatory” knowledge and preserving temporal realism. Second, we extracted a Global Plan from each original case and used it to pre-generate detailed, session-level agendas that guided the AI’s therapeutic focus, maintaining goal continuity and clinical plausibility across long dialogues. Finally, qualified psychology researchers conducted random audits of the generated dialogues, evaluating clinical logic, role adherence, and annotation accuracy; any deficient segments were refined through iterative regeneration or manual correction until they met professional standards—ensuring the dataset’s clinical validity and pedagogical robustness.
D.3 Authoritative Psychology Resources
A substantial portion of our case reports are derived from two distinguished, peer-reviewed Chinese psychology journals, ensuring that the data adheres to rigorous ethical standards and scientific scrutiny: Psychological Monthly***https://www.xlykzz.com/CN/home and Theory and Practice of Psychological Counselinghttps://www.sciscanpub.com/journals/tppc. In addition to empirical case reports, psychology textbooks played a crucial role in our work. They not only helped us extract meta-level therapeutic skills but also served as foundational references for prompt design. Key texts include The System of Counseling and Therapieshttps://book.douban.com/subject/5979615/, Theory and Practice of Counseling & Psychotherapy (Eighth Edition)§§§https://book.douban.com/subject/4179683/, Cognitive Behavior Therapy: Basics and Beyond (Third Edition)https://book.douban.com/subject/36815051/, Existential-Humanistic Therapyhttps://book.douban.com/subject/26304954/, Behavioral Treatment******https://book.douban.com/subject/20494848/, Psychodynamic Formulation††††††https://book.douban.com/subject/26327172/, More Than Miracles——The State of the Art of Solution-Focused Brief Therapy‡‡‡‡‡‡https://book.douban.com/subject/26761621/.
@@ Appendix E Prompts
E.1 Prompt for Client Profile
@@
E.2 Prompt for Therapeutic Plan
@@
E.3 Prompt for Dialogue Features
@@
E.4 Prompt for Session Objective
@@
E.5 Prompt for Skill Suggestion
@@
E.6 Prompt for Dialogue Generation
@@
E.7 Prompt for Client Memory Extraction
@@
E.8 Prompt for Client Merge
@@
E.9 Prompt for Dialogue Summary